# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v4

@@ III.4 New exactly solvable mosaic models
With the universal results presented above, we put forward several new exactly solvable models derived from the type-II QP mosaic model (Zhou et al., 2023) in Eq. (20) and Eq. (22) by removing the MEs. Through a combination of dual transformation, we can construct new nontrivial models, which demonstrate the applications of the above results, and investigate the phase diagram of the models both analytically and numerically. We numerically solve the spectrum and employ the fractal dimension (FD) to phenomenologically characterize the localization properties of the eigenstates $\lvert\Psi\rangle=\sum_{j=1}^{L}u_{j}a_{j}^{\dagger}\lvert\mathrm{vac}\rangle$, where $\mathrm{FD}=-\lim_{L\rightarrow\infty}\ln\sum_{j=1}^{L}|u_{j}|^{4}/\ln L$, with $u_{j}$ being the wave function coefficients, $L$ the system size. In 1D, the FD approaches $1$ for extended states and 00 for localized states, while for critical states, $10<\mathrm{FD}<$ 1. The FD quantifies the effective dimension experienced by an eigenstate: extended states uniformly spread across the system, yielding $\mathrm{FD}=1$. Localized states decay exponentially, effectively perceiving zero dimension, hence $\mathrm{FD}=0$. Critical states, however, exhibit self-similar wave function structures that span the entire system, resulting in an effective dimension between 00 and $1$.
@@
where $\lambda$ is the uniform hopping strengths and $t$ is the strength of balanced QP potential and exchange coupling. The spectrum comprises critical (localized) states for the energies satisfy $E<|\lambda|$ ($E>|\lambda|$(Zhou et al., 2023). Next, we construct new exactly solvable models that host pure critical and localized phases by eliminating the MEs, as outlined in theorem I (Sec. III.1). This is accomplished by removing the spin-independent component of the on-site matrix of the type-II QP mosaic model. This results in a quasiperiodic spin-flipped (QPSF) model given by
@@
This model is exactly solvable, as the hopping coupling matrix satisfies $\det|\sigma_{\pm}|=0$. Applying Avila’s global theory, we find that the system is in the localized phase when QP spin-flipped process dominates $|t|>|\lambda|$, with the analytic localization length $\xi_{l}=2/\log|V/t|$ for all eigenstates. In contrast, when $|t|<|\lambda|$, the system is in the critical phase. The numerical results, as shown in Fig. 3(a), are consistent with the analytical results, with the dashed line marking the phase transition at $\lambda=t$.
@@ IV Exactly solvable models for all fundamental localization phases
From the above subsection we can see that the universal results provide a powerful guidance to construct new exactly solvable models with nontrivial localization physics. In this section, we present the further in-depth study in developing the highly novel models with exact solutions, which enable a comprehensive and unified characterization of all the basic types of MEs and pure phases in quasiperiodic lattices. In particular, we propose two classes of exactly solvable models. The first is 1D spin-selective QP (SSQP) lattice model which is shown to host all basic types of MEs, and the second is the QP optical Raman lattice model which hosts all the seven fundamental localization phases. The seven fundamental phases are the three pure phases (extended, critical, and localized), three coexisting phases for any two of the three types of states, and one coexisting phase with all three types of states.
@@ IV.1 The spin-selective QP lattice model for all fundamental MEs
Figure 4: Generalized incommensurate zeros generated critical states. (a) The quasiperiodic (QP) spin-conserved hopping in this model gives rise to generalized incommensurate zeros (GIZs), which introduce critical orbitals. (b) The spin-selective QP (SSQP) lattice at the exactly solvable regime. Left panel: The model exhibits analytic mobility edges (MEs) at $E_{c}=\pm\lambda^{2}/\mu$, marked by the solid lines, which separate the extended states ($E<|\lambda^{2}/\mu|$) with fractal dimension $\mathrm{FD}\rightarrow 1$, and the rigorous critical states ($E>|\lambda^{2}/\mu|$) with $\mathrm{FD}$ approaching a value between 0 and 1. Right panel: The corresponding $\mathrm{FD}$ for the dual model. The MEs $E_{c}=\pm\lambda^{2}/\mu$ (solid lines) now separate the localized states ($E<|\lambda^{2}/\mu|$) and the critical states ($E>|\lambda^{2}/\mu|$). (c) The SSQP model at another exactly solvable regime that realizes all basic types of MEs. Upper panel: The energy-dependent effective potential resulted from the incommensurately distributed zeros in the shared component, which gives incommensurately distributed divergent potential. Lower panel: The $\mathrm{FD}$ of the eigenstates shown versus $V_{B}/t$ and $E/t$, with $\lambda_{0}/t=1$ and $V_{0}/t=1.5$. The MEs are indicated by the solid lines. All systems have a size of $L=2586$.
@@
$\Pi_{j}=\mu V_{j}^{\mathrm{od}}\Lambda_{-}+\lambda\sigma_{x},\ M_{j}=\lambda\sigma_{x}.$ (30)
@@
Neglecting the constant term in the potential, this effective model can be viewed as 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) without uniform hopping, where the system is in the localized (critical) phase when the QP potential (QP hopping) dominates (see Appendix A for details). Therefore, the system is in the critical phase for $\mu>\lambda^{2}/|E|$ and in the localized phase for $\mu<\lambda^{2}/|E|$. The numerical results shown in Fig. 4(b) right subfigure are consistent with this analytic study. The MEs, $E_{c}=\pm\lambda^{2}/\mu$, separate the critical states for $|E|<\lambda^{2}/\mu$ from the localized states for $|E|>\lambda^{2}/\mu$. These results can also be obtained by applying Avila’s global theory, yielding the LE
@@
The underlying physics of the emergence of the three types of MEs can be achieved by by reducing the SSQP model in the current regime [Eq. (34)] to an effective 1D spinless model, where the effective energy-dependent on-site potential induced by IDZs in shared components plays a key role
@@
When $|E-V_{0}|\leq 2V_{B}$, the system exhibits an incommensurately distributed divergent potential, leading to the critical states. This explains why, when $V_{B}/t$ is sufficiently large, the system exhibits only critical and localized states. In this regime, the analytic LE is
@@
When $|E-V_{0}|>2V_{B}$, the effective potential is finite, giving extended and localized states. The LE can be analytically obtained as
$\gamma(E)=\max\Bigg{\{}\ln|\frac{\chi_{E}+\sqrt{\chi_{E}^{2}-\chi_{B}^{2}}}{|\chi_{0}|+\sqrt{\chi_{0}^{2}-\chi_{B}^{2}}}\Bigg{|},0\Bigg{\}},$ (37)
with $\chi_{B}=2V_{B}/t$, $\chi_{E}=EV_{B}/t^{2}$ and $\chi_{0}=(E-V_{0})/t$. A finite $\gamma(E)$ gives the localized states with localization length $\xi_{l}=\gamma^{-1}(E)$, and $\gamma(E)=0$ corresponds to the extended states.
@@ IV.2 The model with the seven fundamental localization phases
The emergence of the seven fundamental phases can be analytically predicted by combining the universal results and Avila’s global theory. By applying Theorem III in Sec. III.3, the system is reduced to a spinless 1D QP chain with nearest-neighbor hopping when $\det|\Pi_{j}|=0$, which occurs when $|t_{0}|=|t_{\mathrm{so}}|$. We set $t_{0}=t_{\mathrm{so}}$ without losing generality. The system exhibits chiral symmetry at $\eta=1$, where only pure critical and localized phases exist when $t_{\rm so}=t_{0}$, and all three pure phases are present when $t_{\rm so}\neq t_{0}$. Introducing a spin-independent quasiperiodic potential ($\eta\neq 1$) breaks chiral symmetry, leading to another exactly solvable point at $\eta=1/2$, where MEs emerge within the original critical and localized phases for $t_{\rm so}=t_{0}$, giving rise to the L+C phase. Further, various coexisting phases with combinations of the three types of quantum states emerge for $\eta<1$ when $t_{\mathrm{so}}\neq t_{0}$. Therefore, both numerical and analytical results confirm the existence of the seven fundamental localization phases, establishing this system as a universal quantum platform for exploration of localization physics.
@@
Figure 5: Seven fundamental phases realized in 1D quasiperiodic (QP) optical Raman lattice. (a) Model illustration: The hopping coupling matrix includes both spin-conserved and spin-flipped hopping terms, while the on-site matrix incorporates both spin-dependent quasiperiodic (QP) potential coupled to $\sigma_{z}$ and spin-independent QP potential coupled to $\sigma_{0}$. (b) Phase diagram: The phase diagram, shown as a function of the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$, reveals seven distinct phases. These consist of three pure phases: extended (E), localized (L), and critical (C), as indicated in the diagram. In addition, there are three coexisting two-phase regions: (L+E) for the coexistence of localized and extended states, (L+C) for localized and critical states, and (C+E) for critical and extended states. Moreover, a coexisting three-phase region (L+E+C) appears, where all three phases coexist. The phase diagram is obtained with the parameter set $t_{\mathrm{so}}=0.8t_{0}$. (c) Exact Solvable Points: (c1) When $\eta=0$, the entire spectrum is localized. (c2) For $\eta=0.5$, the spectrum splits into localized and critical states, with the mobility edges marked by solid lines at $E_{c}=\pm 2t_{0}$. States with $|E|<2t_{0}$ are critical, while those with $|E|>2t_{0}$ are localized. (c3) At $\eta=1$, the system exhibits pure localized and critical phases, with the transition point marked by the dashed line at $M_{z}=4t_{0}$. The system is in the localized (critical) phase when $M_{z}>4t_{0}$ ($M_{z}<4t_{0}$). The system size used in these calculations is $L=2586$.
@@ IV.2.1 Pure spin-independent quasiperiodic potential
The emergence of NNN hopping begins to destabilize localization, allowing for the appearance of delocalized states. Although the system is no longer analytically solvable for $t_{0}\neq t_{\rm so}$, the possible phases can still be determined. For sufficiently large (weak) $M_{z}$, the system always enters the localized (extended) phase. On the other hand, for intermediate $M_{z}$ values, the chiral symmetry is explicitly broken by spin-independent QP modulation $M_{z}V_{j}^{\mathrm{d}}\sigma_{0}$, leading to the emergence of mobility edges (MEs) between the extended and localized states. This accounts for extended, localized and L+E phases in Fig. 5(b) at $\eta=0$, which extend to small $\eta$ regime with $\eta<\eta_{c}\approx 0.43$.
@@ IV.2.2 Pure spin-dependent quasiperiodic potential
For pure spin-dependent QP potential, where $\eta=1$, we begin by considering the scenario where $t_{0}=t_{\mathrm{so}}$, which yields the following effective hopping and on-site potential
$t_{j}^{\mathrm{eff}}=\frac{-2t_{0}\Delta_{j}}{E},\ V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}+\Delta_{j}^{2}}{E}.$ (44)
@@
$t_{j}^{\mathrm{eff}}=\frac{t_{-}\Delta_{j+1}-t_{+}\Delta_{j}}{E},\ V_{j}^{\mathrm{eff}}=\frac{2t_{0}^{2}+2t_{\mathrm{so}}^{2}+\Delta_{j}^{2}}{E},$ (46)
@@
The system exhibits the chiral symmetry at $\eta=1$, leading to the vanishing MEs and the presence of only three pure phases: localized, extended, and critical. The global theory does not strictly apply here, however, since the system has no mobility edges, the phase boundaries can be determined by analyzing the typical eigenstates. For example, by considering the zero-energy states $E=0$, the eigen-equation Eq. (38) for spin-up (spin-down) states $\psi_{j}$ ($\varphi_{j}$) becomes
$-t_{0}(\psi_{j+1}+\psi_{j-1})-t_{\mathrm{so}}(\psi_{j+1}-\psi_{j-1})+\Delta_{j}\psi_{j}=0,$ (48)
with the $\varphi$ obtained by replacing $t_{\rm so}\to-t_{\rm so}$. The above equation can be expressed in the form of transfer matrix as $(\psi_{j+1},\psi_{j})^{\intercal}=A_{j}(\psi_{j},\psi_{j-1})^{\intercal}$, with $A_{j}$ being
$A_{j}=\left(\begin{array}[]{cc}\Delta_{j}/{t_{+}}&-t_{-}/t_{+}\\ 1&0\end{array}\right),$ (49)
with an analogous form for spin-down states. In the pure spin-dependent limit ($\eta=1$), the delocalized phase is topological while the localized phase is topologically trivial Wang et al. (2020b). The phase boundary between the critical and localized phases can thus be determined by the topological transition. If both eigenvalues of the transfer matrix $A=\prod_{j=1}^{L}A_{j}$ are either less than 1 or greater than 1, the system is topological, and the $\psi$ zero mode is localized at one end of the chain. Assuming $t_{0}>0$ and $t_{\mathrm{so}}>0$, then the two eigenvalues of $A$ satisfy $|\lambda_{1}\lambda_{2}|<1$ , the topological nature is dictated by the larger eigenvalue $|\lambda_{2}|$. After performing a standard similarity transformation DeGottardi et al. (2011, 2013) or applying the global theory, we find the localized-to-critical transition point is given by $|M_{z}/2t_{+}|=1$. By applying the same reasoning to the transfer matrix in the dual space, the phase boundary between localized and critical phases in the dual space is $|M_{z}/2t_{-}|=1$, corresponding to the transition between extended and critical phases in the original space. Therefore, for $\eta=1$ and $t_{0}\neq t_{\mathrm{so}}$, the system exhibits three distinct pure phases as summarized in Table. 2.
@@
Table 2: The criteria for the phases of 1D QP optical Raman lattice model at $\eta=1$, based on the relationship between the QP Zeeman potential $M_{z}$ and the imbalanced hopping $t_{\pm}=t_{0}\pm t_{\mathrm{so}}$.

### v5

@@ III.4 New exactly solvable mosaic models
With the key results presented above, we put forward several new exactly solvable models derived from the type-II QP mosaic model (Zhou et al., 2023) in Eq. (17) and Eq. (19) by removing the MEs. Through a combination of dual transformation, we can construct new nontrivial models, which demonstrate the applications of the above results, and investigate the phase diagram of the models both analytically and numerically. We numerically solve the spectrum and employ the fractal dimension (FD) to phenomenologically characterize the localization properties of the eigenstates $\lvert\Psi\rangle=\sum_{j=1}^{L}u_{j}a_{j}^{\dagger}\lvert\mathrm{vac}\rangle$, where $\mathrm{FD}=-\lim_{L\rightarrow\infty}\ln\sum_{j=1}^{L}|u_{j}|^{4}/\ln L$, with $u_{j}$ being the wave function coefficients, $L$ the system size. In 1D, the FD approaches $1$ for extended states and 0 for localized states, while for critical states, $0<\mathrm{FD}<1$. The FD quantifies the effective dimension experienced by an eigenstate: extended states uniformly spread across the system, yielding $\mathrm{FD}=1$. Localized states decay exponentially, effectively perceiving zero dimension, hence $\mathrm{FD}=0$. Critical states, however, exhibit self-similar wave function structures that span the entire system, resulting in multifractality with FD between 0 and $1$.
@@
where $\lambda$ is the uniform hopping strengths and $t$ is the strength of balanced QP potential and exchange coupling. The spectrum comprises critical (localized) states for the energies satisfy $E<|\lambda|$ ($E>|\lambda|$) (Zhou et al., 2023). Next, we construct new exactly solvable models that host pure critical and localized phases by eliminating the MEs, as outlined in theorem I (Sec. III.1). This is accomplished by removing the $\sigma_{0}$-component of the on-site matrix of the type-II QP mosaic model. This results in a QP spin-flipped (QPSF) model with chiral symmetry as
@@
This model is exactly solvable, as the hopping coupling matrix satisfies $\det|\sigma_{\pm}|=0$. Applying Avila’s global theory (see Appendix E for details), we find that the system is in the localized phase when QP spin-flipped process dominates $|t|>|\lambda|$, with the analytic localization length $\xi_{l}=2/\log|V/t|$ for all eigenstates. In contrast, when $|t|<|\lambda|$, the system is in the critical phase. The numerical results, as shown in Fig. 3(a), are consistent with the analytical results, with the dashed line marking the phase transition at $\lambda=t$.
@@ IV Exactly solvable models for all fundamental localization phases
From the above subsection we can see that the universal results provide a powerful guidance to construct new exactly solvable models with nontrivial localization physics. In this section, we present the further in-depth study in developing the highly novel models with exact solutions, which enable a comprehensive and unified characterization of all the basic types of MEs and pure phases in quasiperiodic lattices. In particular, we propose two classes of exactly solvable models. The first is 1D spin-selective QP (SSQP) lattice model which is shown to host all basic types of MEs, and the second is the QP optical Raman lattice model which hosts all the seven fundamental localization phases. The seven fundamental phases are the three pure phases (extended, critical, and localized), three coexisting phases for any two of the three types of states, and one coexisting phase with all the three types of states. These phases are fundamental, since any generic phase diagram of a QP system can be a certain complex combination of the seven phases.
@@ IV.1 The spin-selective QP lattice model for all fundamental MEs
Figure 4: Generalized incommensurate zeros generated critical states. (a) The quasiperiodic (QP) spin-conserved hopping in this model gives rise to generalized incommensurate zeros (GIZs), which introduce critical orbitals. (b) The spin-selective QP (SSQP) lattice at the exactly solvable regime. Left panel: The model exhibits analytic mobility edges (MEs) at $E_{c}=\pm\lambda^{2}/\mu$, marked by the solid lines, which separate the extended states ($E<|\lambda^{2}/\mu|$) with fractal dimension $\mathrm{FD}\rightarrow 1$, and the rigorous critical states ($E>|\lambda^{2}/\mu|$) with $\mathrm{FD}$ approaching a value between 0 and 1. Right panel: The corresponding $\mathrm{FD}$ for the dual model. The MEs $E_{c}=\pm\lambda^{2}/\mu$ (solid lines) now separate the localized states ($E<|\lambda^{2}/\mu|$) and the critical states ($E>|\lambda^{2}/\mu|$). (c) The SSQP model at another exactly solvable regime that realizes all basic types of MEs. Upper panel: The energy-dependent effective potential resulted from the incommensurately distributed zeros in the shared component, which gives incommensurately distributed divergent potential. Lower panel: The $\mathrm{FD}$ of the eigenstates shown versus $V_{B}/t$ and $E/t$, with $\lambda_{0}/t=1$ and $V_{0}/t=1.5$. The MEs are indicated by the solid lines. All systems have a size of $L=2584$.
@@
$\Pi_{j}=\mu V_{j}^{\mathrm{od}}\Lambda_{-}+\lambda\sigma_{+},\ M_{j}=\lambda\sigma_{x}.$ (27)
@@
Neglecting the constant term in the potential, this effective model can be viewed as 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) without uniform hopping, where the system is in the localized (critical) phase when the QP potential (QP hopping) dominates (see Appendix A for details). Therefore, the system is in the critical phase for $\mu>\lambda^{2}/|E|$ and in the localized phase for $\mu<\lambda^{2}/|E|$. The numerical results shown in Fig. 4(b) right subfigure are consistent with this analytic study. The MEs, $E_{c}=\pm\lambda^{2}/\mu$, separate the critical states for $|E|<\lambda^{2}/\mu$ from the localized states for $|E|>\lambda^{2}/\mu$. These results can also be obtained by applying Avila’s global theory (see Appendix E for details), yielding
@@
The underlying physics of the emergence of the three types of MEs can be achieved by reducing the SSQP model in the current regime [Eq. (31)] to an effective 1D spinless model, where the effective energy-dependent on-site potential induced by IDZs in shared components plays a key role
@@
When $|E-V_{0}|\leq 2V_{B}$, the system exhibits an incommensurately distributed divergent potential, leading to the critical states. This explains why, when $V_{B}/t$ is sufficiently large, the system exhibits only critical and localized states. In this regime, the analytic LE cab be obtained as (see Appendix E for details)
@@
When $|E-V_{0}|>2V_{B}$, the effective potential is finite, giving extended and localized states. The LE can be analytically obtained as (see Appendix E for details)
$\gamma(E)=\max\Big\{\ln|\frac{|\chi_{E}|+\sqrt{\chi_{E}^{2}-\chi_{B}^{2}}}{|\chi_{0}|+\sqrt{\chi_{0}^{2}-\chi_{B}^{2}}}\Big|,0\Big\},$ (34)
with $\chi_{B}=2V_{B}/t$, $\chi_{E}=EV_{B}/t^{2}$ and $\chi_{0}=(E-V_{0})/t$. A finite $\gamma(E)$ gives the localized states, and $\gamma(E)=0$ corresponds to the extended states.
@@ IV.2 The model with the seven fundamental localization phases
The emergence of the seven fundamental phases can be analytically predicted by combining the universal results and Avila’s global theory. By applying Theorem III in Sec. III.3, the system is reduced to a spinless 1D QP chain with nearest-neighbor hopping when $\det|\Pi_{j}|=0$, which occurs when $|t_{0}|=|t_{\mathrm{so}}|$. We set $t_{0}=t_{\mathrm{so}}$ without losing generality. The system exhibits chiral symmetry at $\eta=1$, where only pure critical and localized phases exist when $t_{\rm so}=t_{0}$, and all three pure phases are present when $t_{\rm so}\neq t_{0}$. Introducing a spin-independent quasiperiodic potential ($\eta\neq 1$) breaks chiral symmetry, leading to another exactly solvable point at $\eta=1/2$, where MEs emerge within the original critical and localized phases for $t_{\rm so}=t_{0}$, giving rise to the L+C phase. Away from these exactly solvable limits, the remaining coexisting phases are obtained by continuously deforming the analytically established phase boundaries and are further supported by numerical calculations. Further, various coexisting phases with combinations of the three types of quantum states emerge for $\eta<1$ when $t_{\mathrm{so}}\neq t_{0}$. Therefore, both numerical and analytical results confirm the existence of the seven fundamental localization phases, establishing this system as a novel quantum platform for exploration of localization physics.
@@
Figure 5: Seven fundamental phases realized in 1D quasiperiodic (QP) optical Raman lattice. (a) Model illustration: The hopping coupling matrix includes both spin-conserved and spin-flipped hopping terms, while the on-site matrix incorporates both spin-dependent ($\sigma_{z}$) and spin-independent ($\sigma_{0}$) QP potential. (b) Phase diagram: The phase diagram, shown as a function of the QP Zeeman potential $M_{z}$ and the chiral parameter $\eta$, reveals seven distinct phases: three pure phases–extended (E), localized (L), and critical (C)–and four coexistence regions. The latter include three two-coexistence regimes, (L+E), (L+C), and (C+E), where two types of eigenstates coexist at different energies, as well as a three-phase coexistence region (L+E+C). The phase diagram is obtained with the parameter set $t_{\mathrm{so}}=0.8t_{0}$. (c) Exact Solvable Points: (c1) When $\eta=0$, the entire spectrum is localized. (c2) For $\eta=0.5$, the spectrum splits into localized and critical states, with the mobility edges marked by solid lines at $E_{c}=\pm 2t_{0}$. States with $|E|<2t_{0}$ are critical, while those with $|E|>2t_{0}$ are localized. (c3) At $\eta=1$, the system exhibits pure localized and critical phases, with the transition point marked by the dashed line at $M_{z}=4t_{0}$. The system is in the localized (critical) phase when $M_{z}>4t_{0}$ ($M_{z}<4t_{0}$). The system size used in these calculations is $L=2584$.
@@ IV.2.1 Pure spin-independent quasiperiodic potential
The emergence of NNN hopping begins to destabilize localization, allowing for the appearance of delocalized states. Although the system is no longer analytically solvable for $t_{0}\neq t_{\rm so}$, the possible phases can still be determined. For sufficiently large (weak) $M_{z}$, the system always enters localized (extended) phase. On the other hand, for intermediate $M_{z}$ values, the chiral symmetry is explicitly broken by spin-independent QP modulation $M_{z}V_{j}^{\mathrm{d}}\sigma_{0}$, leading to the emergence of MEs between the extended and localized states. This accounts for extended, localized and L+E phases in Fig. 5(b) at $\eta=0$, which extend to small $\eta$ regime with $\eta<\eta_{c}\approx 0.43$.
@@ IV.2.2 Pure spin-dependent quasiperiodic potential
For pure spin-dependent QP potential with $\eta=1$, we begin by considering the case $t_{0}=t_{\mathrm{so}}$, which yields the following effective hopping and on-site potential
$t_{j}^{\mathrm{eff}}=-2t_{0}\Delta_{j},\ V_{j}^{\mathrm{eff}}=4t_{0}^{2}+\Delta_{j}^{2}.$ (41)
@@
$t_{j}^{\mathrm{eff}}={t_{-}\Delta_{j+1}-t_{+}\Delta_{j}},\ V_{j}^{\mathrm{eff}}={2t_{0}^{2}+2t_{\mathrm{so}}^{2}+\Delta_{j}^{2}},$ (43)
@@
At $\eta=1$, chiral symmetry removes the MEs, leaving only three pure phases: localized, extended, and critical. The global theory does not strictly apply here, however, since the system has no MEs, the phase boundaries can still be determined by analyzing the typical eigenstates. For example, by considering the zero-energy states $E=0$, the eigen-equation Eq. (35) for spin-up (spin-down) states $\psi_{j}$ ($\varphi_{j}$) can be expressed in the form of transfer matrix as $(\psi_{j+1},\psi_{j})^{\intercal}=A_{j}(\psi_{j},\psi_{j-1})^{\intercal}$, with $A_{j}$ the transfer matrix at site $j$. In the pure spin-dependent limit ($\eta=1$), the delocalized phase is topological while the localized phase is topologically trivial Wang et al. (2020b). The phase boundary between the critical and localized phases can thus be determined by the topological transition. If both eigenvalues of the transfer matrix $A=\prod_{j=1}^{L}A_{j}$ are either less than 1 or greater than 1, the system is topological, and the $\psi$ zero mode is localized at one end of the chain. Assuming $t_{0}>0$ and $t_{\mathrm{so}}>0$, then the two eigenvalues of $A$ satisfy $|\lambda_{1}\lambda_{2}|<1$ , the topology is dictated by the larger eigenvalue $|\lambda_{2}|$. By a similar transformation DeGottardi et al. (2011, 2013) or applying the global theory, we find the localized-to-critical transition point is given by $|M_{z}/2t_{+}|=1$. By applying the same reasoning to the transfer matrix in the dual space, the phase boundary between localized and critical phases in the dual space is $|M_{z}/2t_{-}|=1$, corresponding to the transition between extended and critical phases in the original space. Therefore, for $\eta=1$ and $t_{0}\neq t_{\mathrm{so}}$, the system hosts three distinct pure phases as summarized in Table. 2.
@@
Table 2: Phases of 1D QP optical Raman lattice model at $\eta=1$, as a function of QP Zeeman potential strength $M_{z}$ and $t_{\pm}=t_{0}\pm t_{\mathrm{so}}$.